A 750nm light beam in air hits the flat surface of a certain liquid, and the beam is split into a reflected ray and a refracted ray. If the reflected ray is completely polarized when it is at 36.0 degrees with respect to the surface, what is the wavelength of the refracted ray?


1
Expert's answer
2021-03-04T13:15:32-0500

With a grazing angle of "36.0^\u2218" (measured from the surface), the angle of incidence is "54.0^\u2218" , which equals the polarizing angle:

"tan\theta_p = \frac{n_2}{n_1} = \frac{n}{1.00} = n = tan54.0^o = 1.38"

In the liquid, "\lambda_n = \frac{\lambda}{n} = \frac{750nm}{1.38} = 545nm"